OpenStudy (anonymous):

simplify: 5x^3+10x^2+15x / 5x A. x^2+10x^2+15x B. 2x+6 C. x^2+2x+3 D. x^4+2x^3+3x^2

OpenStudy (kamille):

When simplifying, you need to find the same numbers and "letters". So, do you notice any same letters/numbers? Do not forget, that 10=5*2, 15=5*3